The following are demand and supply functions

Qd=17-2P ;Qs=2+3P

Where Qd is quantity demanded,Qs is quantity supplied,and P is price.

I. Find the equilibrium price and quantity demanded and supplied.

II. Given another demand function: Qd=20-3P and also using the demand function in (i), find the market demand when price,P is GH¢2.00.


1
Expert's answer
2021-03-26T11:52:30-0400

Solution:

I.). At equilibrium: Qd = Qs

17-2P = 2+3P

17-2 = 3P+2P

15 = 5P

P = "\frac{15}{5}" = 3


Q = 17-2(3) = 17-6 = 11

Equilibrium price = 3

Quantity demanded and supplied = 11


II). Market demand (Q) = Demand function 1 (Q1) + Demand function 2 (Q2)

Q = 17-2P + 20-3P

Q = 17+20 – 2P-3P

Q = 37 – 5P

Market demand (Q) = 37 – 5P